# Pushing LLMs to Their Logical Reasoning Bound: The Role of Data Reasoning Intensity

### v3

@@ 4 Experimental Settings
    RQ4: How does DRI influence logical reasoning performance in reinforcement learning settings?
4.1 Datasets and Settings
We conduct experiments on four logical reasoning benchmarks: Reclor (contextual reasoning) (Yu et al., 2020), LogicBench (deductive reasoning) (Parmar et al., 2024), LogiQA (workplace logic analysis) (Liu et al., 2020), and LogiQA2.0 (enhanced adversarial patterns) (Liu et al., 2023). For our experiments, we utilize two backbone models: LLaMA3.1-8B-Instruction (Grattafiori et al., 2024) and Qwen2.5-7B-Instruction (Qwen et al., 2025). We refer to them hereafter as LLaMA3.1-8B and Qwen2.5-7B. Additionally, we include GPT-4 (OpenAI, 2023) and DeepSeek-V3 (DeepSeek-AI, 2024) as reference models. For more implementation details, please refer to the Appendix B.
4.2 Analysis of Data Reasoning Intensity (RQ1)
4.2.1 DRI Score Distribution
As shown in Figure 3(a), when the training set is binned into 20 score intervals, the sample counts form a bell-shaped curve with a weighted mean of $\mu=0.526$ and standard deviation $\sigma=0.204$. This approximately Gaussian frequency distribution demonstrates that our DRI scores effectively separate examples by their underlying reasoning potential. Figure 3(a) also plots model error rate against these intervals: error climbs from 8.5% at a score of 0.1 to 59.3% at 0.6, demonstrating a clear positive correlation between DRI and failure rate. Beyond a score of 0.6, error rate plateaus at approximately 61% ± 3.6%. This saturation likely arises because (1) very high DRI items exceed the model’s current reasoning capacity, and (2) the top bins contain few, uniformly hard cases, so failures become uniformly pervasive. Unlike surface metrics (e.g., sentence length or vocabulary complexity), our score distribution captures deeper logical structure.
4.2.2 DRI Score Distribution Balancing
The original test set contains 4,743 examples. Its DRI score distribution (Figure 3(b)) reveals that 74.3% of samples fall into the mid-range interval (0.2, 0.7), while only 17.2% lie in the high DRI interval (>0.7). Such skew can introduce two evaluation biases: (1) Apparent overfitting in the mid-range: strong results in the overrepresented middle interval may conceal weaknesses on higher DRI examples; (2) Insufficient statistical power: the small number of high DRI samples ($n=818$) yields wide confidence intervals for any observed gains. To correct this, we built a balanced benchmark (Figure 3(c)) by drawing 80 examples from each interval: two “extreme” bins (<0.2, >0.9) and contiguous 0.05-wide bins across (0.2, 0.9). This uniform sampling ensures equal representation across the DRI spectrum, eliminating bias and enabling reliable comparisons.
4.2.3 Data Potential Analysis by DRI
Building on our validation of the DRI score and the creation of a balanced test benchmark, we next investigated how filtering training examples by their DRI scores affects learning. As shown in Figure 4, we systematically evaluate model performance when trained on subsets defined by DRI intervals, using both the original and balanced test sets. These experiments reveal three critical patterns, which can be summarized as the following observations:
@@ 4.2.3 Data Potential Analysis by DRI
    Obs 1: Low DRI data can be safely pruned. Both Figure 4(a) and (b) show that training on Range(0.2, 1.0), which omits only the lowest 20% of examples, consistently outperforms full-data training, reducing error rates in nearly every bin. Even when further restricting to Range(0.2, 0.8) (removing both the lowest and highest 20%), overall accuracy remains on par with or slightly above the full-data baseline ($\Delta=+0.3\%$), and mid-range bins (0.2, 0.7) see notable gains. These results confirm that low DRI examples can be pruned without harming model performance and can often lead to improvements, while indiscriminate use of all data may introduce noise.
@@
    Obs 2: High DRI data are catalysts for improvement. Figure 4(a) shows that training on Range(0.2, 1.0) yields the lowest error rates across all bins. When the upper bound is reduced to Range(0.2, 0.9), error rates rise, particularly in the highest DRI bins. Narrowing further to Range(0.2, 0.8) causes a further increase in errors. This stepwise degradation, also reflected in Figure 4(b), confirms that examples with the highest DRI scores drive the most significant performance gains and act as catalysts for model learning.
  • $\bullet$
@@
    Obs 3: Too little data breaks the learning process. Figures 4(a) and (b) show that restricting training to narrow DRI intervals Range(0.3, 0.7) or Range(0.4, 0.7) leads to severe performance degradation, especially in the mid DRI range (0.2, 0.7), where error rates exceed those from broader training ranges. This demonstrates that sparse coverage of the DRI spectrum impairs the model’s ability to internalize and apply core reasoning patterns. The sharp drop in performance underscores the necessity of preserving sufficient data diversity and quantity across all DRI levels to sustain learning.
Figure 4: Experimental results of fine-tuning models in different intervals. "Direct-All" denotes a model fine-tuned on all training examples. "Range(x, y)" denotes a model fine-tuned on examples whose DRI scores fall between $x$ and $y$. Left: Testing set experiment results. Right: Balanced testing set experiment results. The horizontal axis represents score intervals, and the vertical axis represents error rates (lower error indicates better performance).
In summary, these results indicate that training data holds untapped potential and that model’s logical reasoning performance is far from its ceiling. They also reveal limitations of static interval filtering: fixed cutoff thresholds can discard valuable examples or preserve noise. To address these issues, we propose the re-cognizing optimization framework, which first recalibrates the model’s logical reasoning schema and then leverages DRI scores to reinforce its logical reasoning pathways, thereby unlocking the training data’s potential and driving continuous improvements in the model’s logical reasoning performance.
4.3 Re-Cognizing Optimization (RQ2)
Drawing on classic theories of human learning and resource allocation, we propose a two-phase re-cognizing optimization strategy guided by DRI scores. First, Sweller’s cognitive load optimization principle (Sweller, 1988) suggests structuring learning so that basic schemas are established before tackling harder tasks, minimizing extraneous load. Second, the resource-rational analysis (Lieder & Griffiths, 2020) shows that humans allocate effort proportional to task demands, achieving an optimal balance of effort and reward. We leverage these insights as follows.
4.3.1 Phase I: Model Cognition Reshaping
@@
In this phase, we reorder the training data according to DRI scores to "reset" and align the model’s reasoning framework, applying Sweller’s cognitive load theory. Allowing the model to explore the full spectrum of DRI examples from the outset helps it form broad reasoning patterns. This “low‐stakes exploration” mirrors how human learners build foundational knowledge before tackling more challenging tasks, minimizing extraneous cognitive load and establishing a robust framework for subsequent learning.
4.3.2 Phase II: Cognitive Reasoning Enhancement
@@
Here, we implement resource-rational analysis by guiding the model’s focus according to normalized DRI scores. The probability $p$ that the model attends to sample $i$ is calculated as follows:
$p_{i}\;=\;\frac{\hat{s}_{i}}{\sum_{j=1}^{N}\hat{s}_{j}},\quad\hat{s}_{i}=\frac{s_{i}-s_{\min}}{s_{\max}-s_{\min}},$ (9)
@@
where $s_{i}$ is the raw DRI score of sample $i$, $s_{\min}$ and $s_{\max}$ are the minimum and maximum scores in the dataset, $\hat{s}_{i}$ is the normalized score for sample $i$, and the denominator $\sum_{j=1}^{N}\hat{s}_{j}$ sums these normalized scores over all $N$ samples. The re-cognizing optimization stratege echoes Lieder’s insight that "human cognition allocates limited resources to maximize expected reasoning gains relative to reasoning demands". Through cognition‐driven emphasis, the model is steered toward high DRI examples, thus securing greater learning returns under constrained resources.
Figure 5: Experimental results of fine-tuning models using different methods. "Ours-All": model fine-tuned on the full training set using re-cognizing optimization. "Ours-Range(x, y)": model fine-tuned with re-cognizing optimization on examples whose DRI scores fall between $x$ and $y$. Left: Test dataset results. Right: Balanced testset results. The horizontal axis represents score intervals and the vertical axis represents error rates (lower error indicates better performance).
@@
Re-Cognizing Optimization (Ours) 0.930 0.750 0.835 0.857 0.843 0.922 0.775 0.840 0.865 0.851
Re-Cognizing Optimization w/o Stage1 0.904 0.699 0.791 0.747 0.771 0.918 0.726 0.825 0.763 0.806
Re-Cognizing Optimization w/o Stage2 0.910 0.717 0.796 0.824 0.809 0.918 0.768 0.834 0.859 0.844
Qwen2.5-7B Base 0.466 0.372 0.424 0.629 0.509 0.433 0.435 0.409 0.678 0.490
@@
Re-Cognizing Optimization (Ours) 0.944 0.824 0.845 0.858 0.858 0.948 0.797 0.884 0.919 0.886
Re-Cognizing Optimization w/o Stage1 0.856 0.808 0.817 0.833 0.827 0.833 0.790 0.822 0.894 0.835
Re-Cognizing Optimization w/o Stage2 0.798 0.679 0.767 0.827 0.784 0.839 0.684 0.778 0.872 0.793
@@
4.4 Results and Analysis (RQ3)
We evaluate the effectiveness of the re-cognizing optimization strategy by comparing it with static baselines and other data-centric methods across datasets. Other methods include curriculum learning (Bengio et al., 2009) and bin-based progressive learning (Klamkin et al., 2024). In curriculum learning, training examples are introduced in order of increasing DRI scores; in bin-based progressive learning, data is divided into different DRI bins, and the next DRI bin is introduced only after the model has been trained on the current bin. Table 1 summarizes the performance across datasets, while Figure 5 shows the error rates across DRI bins for the original and balanced test sets. Figure 6 (left) compares our method with the other two methods. From the experimental results, we have the following observations:
  • $\bullet$
    Obs 4: Re-cognizing optimization consistently reduces errors across datasets and DRI bins, leading to comprehensive and robust performance gains. In Figure 5, Figure 6 (left), and Table 1, our method consistently achieves the best performance across datasets and test splits, confirming its ability to generalize dataset-specific reasoning patterns. For example, in the lowest DRI interval (0, 0.2), errors drop from 17.6%/25% to 12.6%/11.2% on the original/balanced test sets, demonstrating reduced overfitting to trivial cases. This trend holds across all DRI bins, resulting in a systematic reduction in errors across the spectrum. On datasets like LogiQA and LogiQA2.0, where baseline accuracies are lower, our method shows a significant improvement over others, demonstrating its adaptability and robustness on diverse datasets.
  • $\bullet$
    Obs 5: Re-cognizing optimization demonstrates resilience in data-scarce settings by maintaining strong performance even with restricted training ranges. In Figure 5, when we apply re-cognizing optimization to restricted training sets—Our-Range(0.2, 0.7) and Our-Range(0.2, 1.0)—we observe a 13.6% error increase when high DRI examples (>0.7) are omitted. However, our method effectively compensates for this loss, maintaining performance comparable to full-data training within the Our-Range(0.2, 0.7) subset ($\Delta=+2.5\%$). Expanding the training range to include higher DRI examples (Our-Range(0.2, 1.0)) consistently maintains lower error rates than direct full-data training, validating the predictive power of our DRI and demonstrating that our method can effectively improve performance even in data-scarce settings.
@@ 4.4 Results and Analysis (RQ3)
To assess the contribution of each training stage, we conduct ablation experiments by removing either Stage 1 or Stage 2 from the full re-cognizing optimization pipeline. As shown in Table 1, both variants lead to substantial performance drops on both LLaMA3.1-8B and Qwen2.5-7B. This confirms that both stages are essential for achieving optimal reasoning performance. In summary, our results show our re-cognizing optimization effectively directs the model to high DRI examples, boosting learning efficiency. It systematically reduces errors across DRI bins (proving improved reasoning) and outperforms traditional methods like curriculum and bin-based learning, demonstrating flexibility and robustness across scenarios.
4.5 Expansion Exploration (RQ4)
Some studies (Chen et al., 2025a; Chu et al., 2025; Chen et al., 2025b) suggest that supervised fine-tuning (SFT) and reinforcement learning (RL) play different roles in stimulating the capabilities of language models. To assess the impact of our DRI scores in an RL setting, we adopt the GRPO algorithm (Shao et al., 2024) within the TRL framework (von Werra et al., 2020) and apply LoRA-based parameter-efficient tuning (Hu et al., 2022). We provide proportionally larger accuracy rewards for samples with higher DRI scores, incentivizing the model to handle more reasoning-intensive cases and internalize richer reasoning patterns. We also enforce a structured output format and reward the model for both format compliance and the quality of its reasoning trace. Figure 6(c) and (d) compare our method against two baselines: the base model without RL and a direct variant with fixed accuracy rewards. In both the original and balanced test sets, our DRI-guided method consistently outperforms these baselines, confirming that dynamic, score-based rewards combined with structured output incentives significantly enhance the model’s reasoning capabilities.
@@ 5 Conclusion
This work introduces a data-centric framework for enhancing LLM reasoning by proposing a novel metric—Data Reasoning Intensity (DRI)—to quantify the inherent logical complexity within training samples. We further present the re-cognizing optimization strategy, optimizing training data to better align with LLMs’ reasoning boundaries. Our framework is validated across multiple reasoning benchmarks, demonstrating consistent improvements over existing methods. Additionally, we show that enhancing DRI benefits both supervised and reinforcement learning settings. We hope this study offers new insights into measuring and activating the untapped reasoning potential of LLMs, and inspires future work on cognitive-level data optimization.
Appendix
Appendix A Limitation, Broader Impact, and Future Work
We acknowledge that although our data reasoning intensity score and re-cognizing optimization method are effective for several tasks in our research, they are far from being perfect. Here, we honestly discuss the limitations, broader impact, and potential avenues for future works.
A.1 Limitation
@@
One limitation of our work is the relatively small scale and variety of models we evaluated—due to budget and time constraints, we focused on only two model families (e.g., LLaMA3.1-8B and Qwen2.5-7B), which may limit the generalizability of our findings to larger or more diverse architectures. Additionally, while our data reasoning intensity score provides a useful proxy for reasoning potential, its formulation could be refined further: the current metrics may not capture all aspects of reasoning complexity or transfer seamlessly to other task domains. Moreover, our logical element extraction relies on LLM-based distillation functions, which can introduce noise or inaccuracies; nevertheless, we found these errors to be minor and within acceptable bounds, having minimal impact on overall metric reliability. Finally, our use of GRPO-based reinforcement learning was exploratory and preliminary; more extensive experiments with alternative reward schemes, longer training runs, and varied model capacities will be necessary to fully assess the robustness and scalability of re-cognizing optimization. We also did not explore individualized learning trajectories emphasized in cognitive science, nor use our reasoning-intensity signal to dynamically switch between System 1 and System 2 modes. Moreover, we restrict our evaluation to fine-tuning existing pretrained checkpoints rather than full-scale pretraining, since retraining multi-billion-parameter models from scratch requires resources beyond our current capabilities.
A.2 Broader Impact

### v4

@@ 4.3 Data Potential Analysis by SLKD
    Obs 1: Low SLKD data can be safely pruned. Both Figure 4(a) and (b) show that training on Range(0.2, 1.0), which omits only the lowest 20% of examples, consistently outperforms full-data training, reducing error rates in nearly every bin. Even when further restricting to Range(0.2, 0.8) (removing both the lowest and highest 20%), overall accuracy remains on par with or slightly above the full-data baseline ($\Delta=+0.3\%$), and mid-range bins (0.2, 0.7) see notable gains. These results confirm that low SLKD examples can be pruned without harming model performance and can often lead to improvements, while indiscriminate use of all data may introduce noise.
@@
    Obs 2: High SLKD data are catalysts for improvement. Figure 4(a) shows that training on Range(0.2, 1.0) yields the lowest error rates across all bins. When the upper bound is reduced to Range(0.2, 0.9), error rates rise, particularly in the highest SLKD bins. Narrowing further to Range(0.2, 0.8) causes a further increase in errors. This stepwise degradation, also reflected in Figure 4(b), confirms that examples with the highest SLKD scores drive the most significant performance gains and act as catalysts for model learning.
@@
    Obs 3: Too little data breaks the learning process. Figures 4(a) and (b) show that restricting training to narrow SLKD intervals Range(0.3, 0.7) or Range(0.4, 0.7) leads to severe performance degradation, especially in the mid SLKD range (0.2, 0.7), where error rates exceed those from broader training ranges. This demonstrates that sparse coverage of the SLKD spectrum impairs the model’s ability to internalize and apply core reasoning patterns. The sharp drop in performance underscores the necessity of preserving sufficient data diversity and quantity across all SLKD levels to sustain learning.
In summary, these results indicate that a large volume of data only provides sparse logical knowledge, while the importance of samples with high SLKD has been overlooked. To address these issues, we propose the re-cognizing optimization method in Section 5.
5 SLKD-based Re-cognizing Method
Inspired by cognitive theories (Sweller, 1988; Lieder and Griffiths, 2020), we propose a two-phase re-cognizing optimization. Our optimization method is guided by SLKD and divided into two phases. We first recalibrates the LLM’s logical reasoning schema and then leverages SLKD scores to reinforce its logical reasoning pathways.
5.1 Phase I: Model Cognition Reshaping
@@
In this phase, we reorder the training data according to SLKD scores to "reset" and align the model’s reasoning framework, applying Sweller’s cognitive load theory. Allowing the model to explore the full spectrum of SLKD examples from the outset helps it form broad reasoning patterns. This “low‐stakes exploration” mirrors how human learners build foundational knowledge before tackling more challenging tasks, minimizing extraneous cognitive load and establishing a robust framework for subsequent learning.
5.2 Phase II: Cognitive Reasoning Enhancement
@@
Here, we implement resource-rational analysis by guiding the model’s focus according to normalized SLKD scores. The probability $p$ that the model attends to sample $i$ is calculated as follows:
$p_{i}\;=\;\frac{\hat{s}_{i}}{\sum_{j=1}^{N}\hat{s}_{j}},\quad\hat{s}_{i}=\frac{s_{i}-s_{\min}}{s_{\max}-s_{\min}}$ (8)
@@
where $s_{i}$ is the raw SLKD score of sample $i$, $s_{\min}$ and $s_{\max}$ are the minimum and maximum scores in the dataset, $\hat{s}_{i}$ is the normalized score for sample $i$, and the denominator $\sum_{j=1}^{N}\hat{s}_{j}$ sums these normalized scores over all $N$ samples. The re-cognizing optimization stratege echoes Lieder’s insight that "human cognition allocates limited resources to maximize expected reasoning gains relative to reasoning demands". Through cognition‐driven emphasis, the model is steered toward high SLKD examples, thus securing greater learning returns under constrained resources.
@@
Optimization w/o Stage1 0.904 0.699 0.791 0.747 0.771 0.918 0.726 0.825 0.763 0.806
Optimization w/o Stage2 0.910 0.717 0.796 0.824 0.809 0.918 0.768 0.834 0.859 0.844
Our Re-Cognizing Method 0.930 0.750 0.835 0.857 0.843 0.922 0.775 0.840 0.865 0.851
LLaMA3.1-8B (relative gain) +0.00% +1.08% +1.71% +0.35% +2.06% +1.54% +0.13% +3.83% +0.12% +3.40%
Base 0.466 0.372 0.424 0.629 0.509 0.433 0.435 0.409 0.678 0.490
@@
Optimization w/o Stage1 0.856 0.808 0.817 0.833 0.827 0.833 0.790 0.822 0.894 0.835
Optimization w/o Stage2 0.798 0.679 0.767 0.827 0.784 0.839 0.684 0.778 0.872 0.793
Our Re-Cognizing Method 0.944 0.824 0.845 0.858 0.858 0.948 0.797 0.884 0.919 0.886
Qwen2.5-7B (relative gain) +6.07% +9.87% +4.84% +0.94% +4.25% +4.06% +7.85% +9.68% +0.00% +6.36%
@@
5.3 Results and Analysis
We evaluate the effectiveness of the re-cognizing optimization strategy by comparing it with static baselines and other data-centric methods across datasets. Other methods include curriculum learning (Bengio et al., 2009) and bin-based progressive learning (Klamkin et al., 2024). In curriculum learning, training examples are introduced in order of increasing SLKD scores; in bin-based progressive learning, data is divided into different SLKD bins, and the next SLKD bin is introduced only after the model has been trained on the current bin. Table 1 summarizes the performance across datasets, while Figure 5 shows the error rates across SLKD bins for the original and balanced test sets. Figure 6 (left) compares our method with the other two methods. From the experimental results, we have the following observations:
  • $\bullet$
    Obs 1: Re-cognizing optimization consistently reduces errors across datasets and SLKD bins, leading to comprehensive and robust performance gains. In Figure 5, Figure 6 (left), and Table 1, our method consistently achieves the best performance across datasets and test splits, confirming its ability to generalize dataset-specific reasoning patterns. For example, in the lowest SLKD interval (0, 0.2), errors drop from 17.6%/25% to 12.6%/11.2% on the original/balanced test sets, demonstrating reduced overfitting to trivial cases. This trend holds across all SLKD bins, resulting in a systematic reduction in errors across the spectrum. On datasets like LogiQA and LogiQA2.0, where baseline accuracies are lower, our method shows a significant improvement over others, demonstrating its adaptability and robustness on diverse datasets.
  • $\bullet$
    Obs 2: Re-cognizing optimization demonstrates resilience in data-scarce settings by maintaining strong performance even with restricted training ranges. In Figure 5, when we apply re-cognizing optimization to restricted training sets—Our-Range(0.2, 0.7) and Our-Range(0.2, 1.0)—we observe a 13.6% error increase when high SLKD examples (>0.7) are omitted. However, our method effectively compensates for this loss, maintaining performance comparable to full-data training within the Our-Range(0.2, 0.7) subset ($\Delta=+2.5\%$). Expanding the training range to include higher SLKD examples (Our-Range(0.2, 1.0)) consistently maintains lower error rates than direct full-data training, validating the predictive power of our SLKD and demonstrating that our method can effectively improve performance even in data-scarce settings.
@@ 5.3 Results and Analysis
To assess the contribution of each training stage, we conduct ablation experiments by removing either Stage 1 or Stage 2 from the full re-cognizing optimization pipeline. As shown in Table 1, both variants lead to substantial performance drops on both LLaMA3.1-8B and Qwen2.5-7B. This confirms that both stages are essential for achieving optimal reasoning performance. In summary, our results show our re-cognizing optimization effectively directs the model to high SLKD examples, boosting learning efficiency. It systematically reduces errors across SLKD bins (proving improved reasoning) and outperforms traditional methods like curriculum and bin-based learning, demonstrating flexibility and robustness across scenarios.
5.4 Expansion Exploration
Some studies (Chen et al., 2025a; Chu et al., 2025; Chen et al., 2025b) suggest that supervised fine-tuning (SFT) and reinforcement learning (RL) play different roles in stimulating the capabilities of language models. To assess the impact of our SLKD scores in an RL setting, we provide proportionally larger accuracy rewards for samples with higher SLKD scores, incentivizing the model to handle more reasoning-intensive cases and internalize richer reasoning patterns. We also enforce a structured output format and reward the model for both format compliance and the quality of its reasoning trace. Figure 6(c) and (d) compare our method against two baselines: the base model without RL and a direct variant with fixed accuracy rewards. In both the original and balanced test sets, our SLKD-guided method consistently outperforms these baselines, confirming that dynamic, score-based rewards combined with structured output incentives significantly enhance the model’s reasoning capabilities.
6 Discussion
Our results resonate with the Densing Law of LLMs (Xiao et al., 2024), which observes exponential growth in model "capability density". By positing that Logical Structure is Knowledge, we extend this intuition to the data dimension. This aligns with the view of language modeling as data compression (Delétang et al., 2024), suggesting that reasoning potential depends less on token volume and more on the concentration of logical primitives—a principle validated by the Phi-series (Abdin et al., 2024). We thus advocate for a shift from volume-based Scaling Laws (Kaplan et al., 2020) toward a structural density-aware paradigm.
Regarding a universal data knowledge density, our analysis suggests that density is likely a relative metric tied to LLMs’ cognitive boundary rather than an absolute constant. The gains from high-SLKD samples highlight a logical sparsity in current corpora, where redundant patterns dilute learning signals (Abbas et al., 2023). Future work should investigate whether excessive logical density triggers distribution shifts or overfitting to formal structures, particularly when interacting with LLMs’ internal mechanisms and inductive biases.
@@ 7 Conclusion
In this paper, we presented a data-centric framework to bolster the reasoning capabilities of LLMs. By introducing Structured Logical Knowledge Density (SLKD), we provide a formal metric to quantify the intrinsic logical complexity of training samples. Furthermore, our proposed re-cognizing optimization strategy effectively aligns training data with the inherent reasoning boundaries of LLMs. Extensive experiments demonstrate that our approach consistently outperforms existing baseline methods. We hope our findings offer a fresh perspective on measuring and unlocking the untapped reasoning potential of LLMs, while inspiring future research into cognitive-level data engineering.
Limitation
@@
One limitation of our work is the relatively small scale and variety of models we evaluated—due to budget and time constraints, we focused on only two model families (e.g., LLaMA3.1-8B and Qwen2.5-7B), which may limit the generalizability of our findings to larger or more diverse architectures. Additionally, while our SLKD score provides a useful proxy for reasoning potential, its formulation could be refined further: the current metrics may not capture all aspects of reasoning complexity or transfer seamlessly to other task domains. Moreover, our logical element extraction relies on LLM-based distillation functions, which can introduce noise or inaccuracies; nevertheless, we found these errors to be minor and within acceptable bounds, having minimal impact on overall metric reliability. Finally, our use of GRPO-based reinforcement learning was exploratory and preliminary; more extensive experiments with alternative reward schemes, longer training runs, and varied model capacities will be necessary to fully assess the robustness and scalability of re-cognizing optimization. We also did not explore individualized learning trajectories emphasized in cognitive science, nor use our reasoning-intensity signal to dynamically switch between System 1 and System 2 modes. Moreover, we restrict our evaluation to fine-tuning existing pretrained checkpoints rather than full-scale pretraining, since retraining multi-billion-parameter models from scratch requires resources beyond our current capabilities.
Appendix A Impact, LLM Usage and Future Work
We acknowledge that although our SLKD score and re-cognizing optimization method are effective for several tasks in our research, they are far from being perfect. Here, we honestly discuss the limitations, broader impact, and potential avenues for future works.
A.1 The Use of Large Language Models
We hereby state that LLMs were deployed purely for translation and linguistic refinement tasks. All research ideas, experimental protocols, execution processes, analytical work, and derived conclusions are the exclusive responsibility of the authors. A comprehensive validation of the manuscript has been performed to guarantee that the employment of LLMs did not introduce any false or misleading information.
A.2 Impact